Given the following weekly demand and forecast figures: Week Demand Forecast 1 27 20 2 21 25 3 39 35 The mean square error is: a) 5 b) 7 c) 10 d) 17 e) none of the above
Answer #1

Mean Square error = 1/n*(X-Xi)2 = 1/3 *((27-20)2 +(25-21)2+(39-35)2) = 1/3*81 = 27

Option e none of these

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well